Additional Q5 A, B, and C are independently assorting Mendelian genes controlling the production of a purple pigment. The altPlease show the work for it acting in a pathway, I'm a very confused by this.
Answer #1

Answer:

AABBCC x aabbcc ----- parents

AaBbCc (purple)--F1

AaBbCc x AaBbCc --- F1xF1

AaxAa= A_(3/4)& aa (1/4)

Bb x Bb = B_(3/4)& bb (1/4)

Cc x Cc = C_(3/4)& cc (1/4)

Purple = A_B_C_ = 3/4*3/4*3/4= 27/64

Red = A_B_cc = 3/4*3/4*1/4=9/64

Orange = A_bbC_ + A_bbcc

A_bbC_ = 3/4*1/4*3/4= 9/64

A_bbcc = 3/4*1/4*1/4= 3/64

Orange= 9/64+3/64= 12/64= 3/16

Yellow = 1- (27/64 + 9/64 + 12/64)= 16/64= 1/4